OpenStudy (anonymous):

What's the value of y? 2x+3y=105 x+2y=65 ( Choices ) 15 25 40 65

OpenStudy (wilsondanielle):

start by isolating y (or x) in one of the equations and substituting it into the other. From there, you should be able to solve for the remaining variable using either equation and yield the value of y.

OpenStudy (wilsondanielle):

eg: x=65-2y 2x+3y = 2(65-2y) + 3y =105 can you find the value of y?

OpenStudy (triciaal):

multiply the second equation by 2 then subtract from equation 1

OpenStudy (anonymous):

i was gonna type what triciaal wrote

OpenStudy (anonymous):

I'm kinda lost.

OpenStudy (anonymous):

oops sorry this is a high school question

OpenStudy (anonymous):

unless you know rref u wouldnt understand it my bad

OpenStudy (anonymous):

Lol yeah xD

OpenStudy (triciaal):